There are 53 oranges in a box that contain 77 total peices of fruit. How many oranges need to be added so the box contains 84% oranges?
Answer by edjones(8007)

x=number of oranges that must be added.
(53+x)/(77+x)=.84
53+x=.84(77+x)
53+x=64.68+.84x
.16x=11.68
x=73
.
Ed